OpenStudy (anonymous):

the square root of 32y to the 7 power

jhonyy9 (jhonyy9):

so squareroot sign exponent 1/2 --- yes ?

jhonyy9 (jhonyy9):

32 is 2 on exponent 5

jhonyy9 (jhonyy9):

and y is on exponent 7

jhonyy9 (jhonyy9):

so (a^b)^c = ???

jhonyy9 (jhonyy9):

so than there is (2^5 *y^7 )^(1/2) - how many will be this ?

OpenStudy (anonymous):

omg i have no idea i dont even know how to read that

jhonyy9 (jhonyy9):

so if there is sqrt(x^a) =(x^a)^(1/2) =x^(a *(1/2)) =x^(a/2) can you using this example in your exercise ?

OpenStudy (anonymous):

they give 4y^3 square root 2y 9 3y square root of 2 2y^3 square root of 2y

jhonyy9 (jhonyy9):

how do you think what will be right ?

jhonyy9 (jhonyy9):

if you replace 4 inside squaroot how many will be ?

jhonyy9 (jhonyy9):

will be 4 squared =4^2 = 4*4 = ?

jhonyy9 (jhonyy9):

so you have wrote : they give ...

jhonyy9 (jhonyy9):

check these answers for you can giving the right answer

jhonyy9 (jhonyy9):

32 =2^5 =2^4 *2 but you know that 2^4 inside squaroot = 2^2 so than sqrt(2^5) = sqrt(2^4 *2) = =2^2 *sqrt2 =4sqrt2 ok ?

jhonyy9 (jhonyy9):

came on

OpenStudy (anonymous):

ooooooh lol okay thanks i got it

jhonyy9 (jhonyy9):

good luck bye